OpenStudy (anonymous):

what is the solution to the equation 5x -2y =-9 2,-4 1,-3 -1,2 -2,5

OpenStudy (kitty626oldrose):

what do u think it is

OpenStudy (greatlife44):

easiest way to do this is to plug in your x and y values

OpenStudy (anonymous):

Idk how to do this can u like explain it to me

OpenStudy (greatlife44):

yeah sure. it's actually like this when we have an equation what we have on the right side must equal what we have on the left.

OpenStudy (anonymous):

ok but..how do we make it equal?

OpenStudy (greatlife44):

so, the correct answer will be the one where your answer gets you 9 when you plug in x and y

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

i got B

OpenStudy (mathmale):

Please express your solutions as points: enclose each set of numbers inside parentheses: 2,-4 1,-3 -1,2 -2,5 becomes (2,-4), (1,-3), and so on. Is (2,-4) a solution? This asks, "when x=2 and y=-4, is 5x -2y =-9 true or false? Please rewrite 5x -2y =-9. Let x=2 and let y=-4. Is the resulting equation true?

OpenStudy (greatlife44):

@study_123 can you show us your work?

OpenStudy (mathmale):

Note: You don't "make it equal." Rather, you CHECK whether the equatino is true for a given point (x,y).

OpenStudy (anonymous):

OK, 5*1-2*-3=-9

OpenStudy (greatlife44):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @study_123 OK, 5*1-2*-3=-9 \(\color{#0cbb34}{\text{End of Quote}}\) don't forget your ( ) 5(1)-2(-3) = -9?

OpenStudy (anonymous):

Yes

OpenStudy (greatlife44):

5(1) = 5 and 2(-3) = - 6 but we have 5-(-6) = 5+6 = 11

OpenStudy (anonymous):

so you're saying my answer is wrong?

OpenStudy (anonymous):

The answer is C.

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok hold up lemme check